Please help with this question. My son needs help. Thank you so much!
Question:
Two rectangles are similar. One is 5 cm by 12 cm. the longer side of the second rectangle is 8 cm greater than twice the shorter side. find its length and width.




Answer by jim_thompson5910(35256)

5/12 = x/(2x+8)


5(2x+8) = 12x


10x+40=12x


10x=12x-40


10x-12x=-40


-2x=-40


x=(-40)/(-2)


x=20

So the shorter side is 20 cm and the longer side is 2*20 + 8 = 48 cm